OpenStudy (anonymous):

f(x) = 2x - 2 g(x) = x + 4 Find (f+g)(x) HELP

OpenStudy (anonymous):

add

OpenStudy (anonymous):

\((f+g)(x)=f(x)+g(x)=2x-2+x+4=...\)

OpenStudy (anonymous):

I know that but I'm having trouble solving it

OpenStudy (anonymous):

I got 2/3

OpenStudy (anonymous):

You shouldn't get a value... In general if you have two functions f(x) and g(x) then their sum h(x) = f(x) + g(x) will be a function of x. In your case you have 2x-x = x so it can't be a constant...